OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = –4 and a5 = –1,024?

OpenStudy (anonymous):

what do you not understand here?

OpenStudy (anonymous):

a(r)^(n-1) -1024/-4 = 256 r^4 = 256 r = 4 (-4)(4)^(n-1) = (-4)(4)^(7-1) = -16384